OpenStudy (3psilon):

3^x = 5^y what is x/y ? Plz help . I'm terrible with Logs

OpenStudy (anonymous):

Bah, logs with different bases . . . Gimme a sec to remember how to do this.

OpenStudy (anonymous):

Ok, got it. Take \(\large log_3\) of both sides.

OpenStudy (anonymous):

Then you can use the logarithm power rule.

OpenStudy (3psilon):

Do you mind working please? @CliffSedge . I'm a little rusty on the rules

OpenStudy (anonymous):

Oh, I suppose so. It's pretty easy.. Given \(\large 3^x = 5^y\) \(\large log_3\) both sides \(\large log_3(3^x) = log_3(5^y)\) \(\large x = 3^x = log_3(5^y)\) Understand so far?

OpenStudy (anonymous):

*oops, never mind that = "3x" in the last line. I thought I deleted that.

OpenStudy (3psilon):

I'm sorta catching on

OpenStudy (anonymous):

Should be \(\large x=log_3(5^y)\)

OpenStudy (anonymous):

Do you see that \(\large log_3(3^x)=x\) ?

OpenStudy (anonymous):

The logarithm function tells you what the exponent is on that base.

OpenStudy (anonymous):

So, since a logarithm is an exponent, and if you have a power-to-a-power, you multiply exponents, \(\large log_3(5^y)=y log_3(5)\) . I think you can get the rest from here.

zepdrix (zepdrix):

 Still having some trouble pimp or you figured this one out already. Figured I'd throw my 2 cents just in case :D Sorry that it's a little bit messy, hopefully you can follow it. The asterisks are log rules.